OpenStudy (anonymous):

Find an equation of the tangent line to y = 2 sin x at x = π/6

OpenStudy (anonymous):

take the derivative, get \[2\cos(x)\] then replace x by \[\frac{\pi}{6}\]

OpenStudy (anonymous):

youre answer is in like slashes and [][ so it looks like \2\xos(x)\] \frac{pi6\]

OpenStudy (anonymous):

so what is the answer to it?

OpenStudy (mertsj):

He said to take the derivative and get 2cos(x). Then replace x with pi over 6

OpenStudy (anonymous):

my homework said 2cos(pi/6) was wrong

OpenStudy (mertsj):

When you take the derivative, you get the slope function. When you replace x with pi over 6 you are finding the slope of the tangent at pi/6. Then you must use that slope to find the equation of the tangent line which the problem requires.

OpenStudy (anonymous):

so is 2cos(pi/6) the answer or does it have more to it?

OpenStudy (mertsj):

What are you supposed to find?

OpenStudy (anonymous):

Find an equation of the tangent line to y = 2 sin x at x = π/6

OpenStudy (mertsj):

Do you know what the equation of a line looks like?

OpenStudy (anonymous):

y = mx+b?

OpenStudy (mertsj):

yes. So I don't think 2cos(pi/6) looks much like the equation of a line, do you?

OpenStudy (anonymous):

nope, pi/6x + 2cos? or the other way around?

OpenStudy (mertsj):

Can you evaluate 2cos(pi/6)?

OpenStudy (anonymous):

i have no idea?

OpenStudy (mertsj):

\[2\cos \frac{\pi}{6}=2\times \frac{\sqrt{3}}{2}=\sqrt{3}\]

OpenStudy (mertsj):

Does that look familiar?

OpenStudy (mertsj):

Have you ever studied the special right triangles or the unit circle?

OpenStudy (anonymous):

it has lines and stuff in it... does it say 2cos(pi/6) = 2 * sqrt3/2

OpenStudy (mertsj):

Yes. And that 2(sqrt3/2)=sqrt3

OpenStudy (aravindg):

could u help me pls after this?

OpenStudy (mertsj):

What do you need help with?

OpenStudy (anonymous):

so in all the equation is?

OpenStudy (mertsj):

So sqrt3 is the slope of the tangent line.

OpenStudy (mertsj):

And it goes through a point on the graph of y = 2sin(x) that has an x coordinate of pi/6

OpenStudy (anonymous):

so sqrt3x+cos(pi/6)

OpenStudy (mertsj):

If pi/6 is the x value, then 1 is the y value. So now we know the slope of the line and a point on the line. We should be able to write the equation using point slope form.

OpenStudy (mertsj):

y-1=sqrt3(x-pi/6)

OpenStudy (mertsj):

And that is the equation of the tangent line.

OpenStudy (anonymous):

this is way too much work for a problem lol, so its y - 1 = sqrt3(x-pi/6)

OpenStudy (mertsj):

It's hardly any work at all if you know what you are doing.

OpenStudy (anonymous):

the homework said it was wrong :\

OpenStudy (anonymous):

y = 2sinx --> f(pi/6) = 2sin(pi/6) = 1 y' = 2cosx At x = pi/6, y' = 2cos(pi/6) = sqrt(3) Tangent line: y - 1 = V3 ( x - pi/6) =>y = V3x - piV3/6 + 1

OpenStudy (anonymous):

Then you should double check, otherwise I'm sure your homework solution's misprinted!

OpenStudy (anonymous):

so is it y = sqrt3x - (pisqrt3/6) + 1

OpenStudy (anonymous):

In short form, it is: y - 1 = V3 ( x - pi/6)

OpenStudy (anonymous):

Most of the time, people want slope y intercept form, so you need to multiply it out.

OpenStudy (anonymous):

so is that how i should put it in? how i did it?

OpenStudy (anonymous):

Since tangent line is straight line: y = mx + b

OpenStudy (anonymous):

you need to find slope m!

OpenStudy (anonymous):

m = derivative of function at specific point!

OpenStudy (anonymous):

If you still have question, just let know!

OpenStudy (anonymous):

i got it wrong

OpenStudy (anonymous):

How, show me!

OpenStudy (anonymous):

idk, i put it in like that and it was my fifth try so it said i was wrong

OpenStudy (anonymous):

idk, i put it in like that and it was my fifth try so it said i was wrong

OpenStudy (anonymous):

Why don't you try to V3 = 1.73, pi = 3/14, calculate yourself My approximation is y = 1.73x + .0931